Fermi National Laboratory
 
Operation Summary for April 11, 2012 - April 13, 2012
 

Wednesday April 11

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, and SeaQuest experiments taking beam.

During the day shift, at 8:15 AM, Preaccelerator personnel accessed the I- Source Dome to make repairs. At 9:37 AM, the NuMI beam permit tripped off for the first of four times in three and a half hours due to a power supply’s (E:V118) output drifting high; everything reset. At 9:41 AM, the NuMI beam permit tripped off due to a high magnet temperature indication on HV101; it reset and the temperature limit was raised slightly. Muon Ring personnel began one-hour study at 2:12 PM. A MI person began a study around 3:12 PM. SeaQuest, which had been in access for most of this shift, resumed taking beam at 3:53 PM.

During the evening shift, a MI kicker (I:KPS5S) tripped off for the first of two times in five minutes; it reset. The MI study ended at 5:38 PM. Operations competed a Booster injection tune at 9:32 PM. Operations established beam to the Meson FTBFT-992 experiment at 10:30 PM. SeaQuest accessed several times during this shift.

Thursday April 12

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-992.

During the midnight shift, a Booster RF station (BRF1) tripped off at 1:56 AM and would not reset; Operators bypassed the station. A Muon Ring person began a study at 4:35 AM. At 5:19 AM, BRF11 tripped off for the first of four times in two hours due to a ground fault indication; it reset. At 6:45 AM, the NuMI beam permit tripped off for the first of two times in one hour due to V118’s output drifting low; everything reset. At 7:54 AM, the NuMI beam permit tripped off due to HV101; it reset. T-992 made several accesses during this shift.

During the day shift, the Muon Ring study ended at 8:35 AM. At 8:36 AM, Linac Preaccelerator personnel switched from the I- Source to the H- Source. RF technicians reported that they had repaired BRF1 at 8:43 AM. A Muon Ring person began a reverse proton study at 10:28 AM. At 10:32 AM, BRF9 tripped off for the first of seven times in two hours; it reset. The NuMI beam permit tripped off at 10:44 AM due to V118; it reset. RF techs investigated the problems with BRF9 at noon. At 12:43 PM, Booster and MI personnel began a phase lock study; this study ended at 3:33 PM. SeaQuest made an access this shift.

During the evening shift, Operations disabled beam to T-992 at 4:17 PM. The Muon Ring reverse proton study ended at 5:36 PM. FTBF experiment T-1005 took over the MTest beam at 8:20 PM. MI personnel completed a tune of the MI-8 transfer line at 7:31 PM. Operations completed a Booster injection tune at 8:40 PM.

Friday April 13

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1005.

During the midnight shift, a Muon Ring person began a study at 4:29 AM. At 7:07 AM, Operations received a report that the LRF1 water skid was making noise; a Water Group person investigated and reported that the cavity pump bearings were going bad and needed to be replaced. T-1005 made an access this shift.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The OpsCo agrees with Booster, MI, and Muon Ring study plans for today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/13/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory